OpenStudy (anonymous):

Solve: -7.2 – (+3.8)

OpenStudy (anonymous):

doesn't anyone own a calculator?

OpenStudy (anonymous):

-4.3

OpenStudy (anonymous):

\(-7.2-3.8=-(7.2+3.8)=-11\)

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=-7.2-3.8

OpenStudy (anonymous):

not sure where the \(4.3\) came from if the signs are both negative, ignore the signs, add, and then stick a minus sign out front for example \(-10-20=-30\)